a single phase overhead transmission line delivers 5000 kw at 11 KV and 0.8 factor lagging. total resistance and inductive reactance per conductor is 0.45 ohm and 0.08 ohm respectively , calculate line losses
1
Expert's answer
2021-08-09T05:34:02-0400

"I = \frac{P}{V_R cos(\phi_R)} = \frac{5000\times 100}{11000\times 0.8}=568.18A"


Line loss, "I^2R = (568.18^2 \times 0.45) = 145.27 KW"